Falconmaster:

Two cards are randomly drawn from an ordinary deck of playing cards without replacement. What is the probability that the first card is a heart and the second card is a diamond?

Falconmaster:

@Hero

dude:

Do you have an idea on how to solve?

Falconmaster:

fractions then calculate the mas of the sun times the speed out sound if a = 6

dude:

LOL Seriously though, are you lost? Maybe the lessons from class should help you (Have you done a tree diagram?)

Falconmaster:

well at first is a fair 25% that it will be hearts then idk the rest

Falconmaster:

tree? um no they never taught me that

Hero:

@Falconmaster so what is the probablity of drawing the first card?

Falconmaster:

25%

Falconmaster:

cause there is 4 suits and 11 cards or something like that in each one

Falconmaster:

oh 13 not 11

Falconmaster:

OH then there is a 33%??? chance of drawling???/ im wrong or would it be another 25%?

Hero:

It asks for the probability of drawing a heart as the first card, which means you should ask yourself the following set of questions: How many hearts are there in total in the deck? Next ask yourself, "How many cards are there in the deck in total?". Then divide

Falconmaster:

waitv 2/52

Falconmaster:

13 then 13 divied by 52?

Hero:

\(\dfrac{\text{Number of hearts in the deck}}{\text{Total number of cards in the deck}}\)

Hero:

Correct

Falconmaster:

0.25

Hero:

Okay so now, the next step says, "without replacement" so that means when you calculate the probability of drawing the second card, you will have one less card in the deck than you had previously. Keep that in mind.

Falconmaster:

13 divied by 51 or 13 divied by 52 again

Hero:

Now ask yourself, "How many diamonds are there in the deck?" and then "How many total cards are there in the deck?". And divide that. Then afterwards multiply the result of the two probabilities

Falconmaster:

13 divied by 52 so .25 times 2?

Falconmaster:

.5?

Hero:

There's one less card in the deck remember? I was only explaining the process. I didn't say anything you already wrote was incorrect.

Falconmaster:

X.X XD lol okay

Hero:

But I definitely would discourage from guessing. Do you know know what the two fractions are you will be mulitplying?

Falconmaster:

i think 13/52 and 13/51 but idk im not good with math

Falconmaster:

im thinking .5 cause i can jk cause of the make but then idk

Falconmaster:

*math

Falconmaster:

@Hero would it be 13/17?

Falconmaster:

or .7?

Hero:

13/52 and 13/51 is correct.

Falconmaster:

ooo so i am smart XD

Falconmaster:

i used math way and divied the fractions and got .9

Hero:

Multiply this: \(\left(\dfrac{13}{52}\right)\left(\dfrac{13}{51}\right)\)

Falconmaster:

HA 0.06

Hero:

There's only one right answer. You're allowed to reduce the answer to a fraction, but not to a decimal. The general rule is, if you start with fractions you should end with fractions.

Hero:

Please do not use a calculator for this.

Hero:

There is no need. You already know 13/52 reduces to 1/4

Falconmaster:

oh whaleden

Hero:

So really, all you're multiplying is \(\left(\dfrac{1}{4}\right)\left(\dfrac{13}{51}\right)\)

Hero:

Also understand that I personally do not like decimals and only resort to them when I absolutely have to which is not often.

Falconmaster:

The when you got your fractions mutiplyed then you can turn it in to a decimal and you should end up with 0.06 ill try that next time

Falconmaster:

oh sorry forgot to add that the answer was a decimal

Falconmaster:

shanks hero

Hero:

Did you understand anything I said about the fact that if you start with fractions you should end with fractions?

Hero:

This is a general rule of mathematics: end with the same notation you started with.

Falconmaster:

Yeah i heard you id prefer to do it that way but stupid study island has to make it rocket science

Hero:

The correct answer is "The probability of drawing a heart and then a diamond is \(\dfrac{13}{204}\).

Falconmaster:

for some dumb reason they wanted a decimal which im like no